# Cascaded Multi-path Shortcut Diffusion Model for Medical Image Translation

### v1

@@ 4 Discussion
The presented work also has limitations with several potential improvements that are important subjects of our future studies. Firstly, we only validated our method on three different modalities, and evaluations on more diverse applications could be included. Even though we framed CMDM as an image-space post-processing tool here, we believe it can be further tailored to specific translation problems. For example, we could include physic-informed modules, such as data consistency [27, 31], in CMDM which may further improve its applications in medical image reconstruction [39, 29]. Secondly, the current CMDM is implemented in a 2D fashion, while 3D is desirable in many medical image translation tasks. Theoretically, we could directly change all the networks in CMDM into 3D networks to enable 3D applications, but it may be infeasible with the current computation power. Alternatively, we could also utilize multi-view diffusion or 2.5D or memory-efficient strategies to scale CMDM into 3D [4, 35, 1]. Thirdly, the inference speed is still considered relatively long as compared to previous methods, especially the classic CNN-based methods. While we discussed the trade-off between performance and speed above, it is also desirable to maintain optimal performance with increased inference speed. Utilizing accelerated diffusion models, such as DDIM and Resshift [30, 38], in CMDM could potentially help achieve this goal. Lastly, CMDM requires data with paired images for training, but such data is not always available. Unpaired translation diffusion model strategies [26, 22] could also potentially be deployed here to mitigate this challenge. Moreover, future works also include evaluations of downstream clinical applications of these translated images.
@@ Acknowledgments
This work was supported by the National Institutes of Health (NIH) grant R01EB025468, grant R01CA275188, and Yale PhD fellowship.

### v2

@@ 5 Discussion
The presented work also has limitations with several potential improvements that are important subjects of our future studies. Firstly, we only validated our method on three different modalities, and evaluations on more diverse applications could be included. Even though we framed CMDM as an image-space post-processing tool here, we believe it can be further tailored to specific translation problems. For example, we could include physic-informed modules, such as data consistency [35, 40], in CMDM which may further improve its applications in medical image reconstruction [49, 37]. Secondly, the current CMDM is implemented in a 2D fashion, while 3D is desirable in many medical image translation tasks. Theoretically, we could directly change all the networks in CMDM into 3D networks to enable 3D applications, but it may be infeasible with the current computation resources. For example, we attempted to employ the 3D CMDM with an input size of $256\times 256\times 128$ on an 80G H100 GPU, however, it cannot fit into the memory even with a single batch size. Alternatively, we could also utilize multi-view diffusion or 2.5D or memory-efficient strategies to scale CMDM into 3D [5, 45, 2] which will be extensively investigated in our future works. Thirdly, the inference speed is still considered relatively long as compared to previous methods, especially the classic CNN-based methods. While we discussed the trade-off between performance and speed in the previous paragraph, it is also desirable to maintain optimal performance with increased inference speed. Utilizing accelerated diffusion models, such as DDIM and Resshift [38, 48], in CMDM could potentially help achieve this goal. To accelerate the inference speed for time-critical clinical scenarios, such as real-time translation in intervention radiology, one could also consider alternative solutions. For example, we could also consider distilling the diffusion model knowledge into the one-step inference GAN model [17], such that GAN with diffusion model performance and real-time capability can be realized. Fourthly, in the current implementation of CMDM, we did not implement ways to monitor the first step’s image generation. If unsatisfactory results were generated in the first step, the error could propagate to the next step. However, this should be reflected on the CMDM final uncertainty map where increased uncertainty value, i.e. pixel-wise standard deviation, should be observed. On the other hand, we could also further include uncertainty estimation techniques, e.g. Monte Carlo Dropout [8], in the first step of cGAN, thus monitoring prior image generation. Lastly, CMDM requires data with paired images for training, but such data may not always be available in certain applications. Unpaired translation diffusion model strategies [34, 30] could also potentially be deployed here to mitigate this challenge. For example, one could consider using CycleGAN to generate the prior image, and then using a multi-path version of UNIT-DDPM [34] for further refinement of the prior image. This is an interesting direction to be investigated in our future works. Moreover, future works also include evaluations of how CMDM impacts the downstream clinical applications. For example, we will investigate if the CMDM-translated images provide similar lesion detection capability or radiomic features when compared to the ground truth images, thus validating the clinical values of our method.
@@ Acknowledgments
This work was supported by the National Institutes of Health (NIH) grant R01EB025468 and grant R01CA275188.